OpenStudy (nsh4267):

The population N(t) (in millions) of a country t years after 1980 may be approximated by the formula N(t) = 205e0.0104t. When will the population be twice what it was in 1980?

OpenStudy (solomonzelman):

The formula is, \(\large\color{black}{N(t) = 205e^{0.0104t} }\) However when will it be twice more than in 1980 depends on what the value of N(t) (or the present population) is.

OpenStudy (nsh4267):

This is all that was in the problem

OpenStudy (solomonzelman):

well, let me show why I said what I said, and what I still hold. \(\large\color{black}{ N(t)=205e^{0.0104t} }\) hitting both sides with a natural log, (the ln) \(\large\color{black}{ \ln(~N(t)~~)=\ln(~205e^{0.0104t} ~) }\) exponent will go in front of the natural log, \(\large\color{black}{ \ln(~N(t)~~)=0.0104t~\ln(~205e ) }\) apply the rule ln(a times b)=ln a + ln b \(\large\color{black}{ \ln(~N(t)~~)=0.0104t~[\ln(~205)+\ln(e )] }\) ln (e) =1 \(\large\color{black}{ \ln(~N(t)~~)=0.0104t~[\ln(~205)+1] }\) and in this equation (just like before), you can't solve for 2 unknown variables. (since it is a single equation with more than 1 unknown). the "t" or number of years when N(t) doubles, would depends on the principal value of N(t), saying what N(t) is NOW.

OpenStudy (solomonzelman):

I don't think it has a solution, although using a natural log is certainly the approach when N(t) is given but the "t" is missing. Sorry that I can't help you.